# compat=1.1.0

### v1

@@ Abstract
The canonical scaling of the Hall angle, $\cot\theta_{H}\propto T^{2}$, observed in optimally doped cuprates and several strange metals, has been widely interpreted as evidence for distinct transport and Hall scattering rates [53]. This behavior, however, is not universal. In electron-doped cuprates, such as $\mathrm{Pr_{2-x}Ce_{x}CuO_{4-\delta}}$ [51], and in layered ruthenates including $\mathrm{Sr_{2}RuO_{4}}$ and $\mathrm{Sr_{3}Ru_{2}O_{7}}$ [52], clear departures from the $T^{2}$ law have been reported. The coexistence of multiple Fermi-surface sheets with distinct orbital characters leads to orbital-dependent scattering lifetimes that evolve differently with temperature and magnetic field. In $\mathrm{Sr_{2}RuO_{4}}$, for instance, the $\alpha$, $\beta$, and $\gamma$ bands derived from different Ru $4d$ orbitals exhibit markedly different coherence scales and gap-like features [52]. Such orbital-selective effects effectively renormalize the temperature dependence of the transport tensors, producing an apparent breakdown of the simple scaling. These observations point to the necessity of introducing a material-dependent cutoff scale $T_{0}$, which encodes band-structure and orbital complexity within a unified transport framework.
@@
$A(\omega;B)\simeq\frac{\kappa_{b}}{\bar{C}}\frac{|B|\,|\omega|}{T_{0}+T},$ (42)
@@
$\Gamma(0)=\frac{\kappa\pi}{8\bar{C}}\frac{|B|\,T^{2}}{T_{0}+T},$ (43)
so that at fixed $T$ the longitudinal resistivity scales linearly with field, $\rho_{xx}\propto|B|$. This mechanism captures key experimental observations in quantum-critical strange metals. In BaFe2(As1-xPx)2, near the QCP the scattering rate follows [54]
$\Gamma\approx\sqrt{(\alpha k_{B}T)^{2}+(\eta\mu_{B}B)^{2}},\quad\frac{\eta}{\alpha}\approx 1,$ (44)
demonstrating nearly identical thermal and magnetic energy scales. Similarly, in overdoped La2-xSrxCuO4, scale-invariant linear-in-$B$ resistivity is observed up to $80~\mathrm{T}$ [55], with slope $\beta\sim\alpha k_{B}/\mu_{B}$, consistent with Eq. (43). Together, these results are consistent with a Planckian-limited scaling,
$\Gamma(T,B)\propto\max\{\alpha k_{B}T,\eta\mu_{B}B\},$ (45)
@@
Acknowledgments- We thank Xiao Hu, Sang-Jin Sin, Yan Liu, Matteo Baggioli and Kui Jin for helpful discussions. We would also like to express our gratitude to colleagues in the "Eight Immortals Crossing the Sea" WeChat group for their engaging discussions on topics beyond physics. This work was partially supported by NSFC, China (Grant Nos. 12275166 and 12311540141).
@@ Appendix A DC Conductivity of Fermi liquid
We present the standard Kubo-bubble derivation and show its reduction to the semi-classical (relaxation-time) form used in the main text.
@@ A.2 DC limit and the quasi-particle approximation
where $\partial_{\epsilon}f(\epsilon)$ restricts contributions to energies within a thermal window $\sim T$ around the Fermi level. Here, $A(\mathbf{k},\epsilon)$ is the single-particle spectral function, and $v_{x}(\mathbf{k})=\partial\varepsilon_{\mathbf{k}}/\partial k_{x}$ is the group velocity along the $x$-direction.
Near the Fermi surface, the spectral function is well approximated by a Lorentzian form
@@
is the single-particle density of states (DOS). Equation (53) is equivalent to the Boltzmann relaxation-time approximation when vertex corrections are negligible, capturing the essential semiclassical physics of momentum relaxation [50].
In a Landau Fermi liquid, the low-energy scattering rate scales as
@@ Appendix B Fermi-liquid limit: Quadratic magnetoresistance
which arises when the magnetic field does not generate new low-frequency scattering channels. This includes impurity- or phonon-dominated elastic scattering (field-independent $\tau_{\rm tr}$) and conventional FL inelastic scattering with $\tau_{\rm tr}^{-1}\propto T^{2}$. The expansion is valid in the weak-field limit $\omega_{c}\tau_{\rm tr}\ll 1$ and continuous spectral densities. Departures from this behavior—such as Landau quantization or field-enhanced soft modes—lead to non-FL magnetoresistance, including the linear-in-$|B|$ response discussed in the main text.
@@
Appendix C Experimental estimation of $T_{0}$
The aim of this section is to determine the emergent Tsallis scale $T_{0}$ that governs non-Fermi liquid scattering in correlated materials. This scale enters the parametrization $q(T)=1+T/T_{0}$ of the Tsallis distribution and sets the strength of soft-mode fluctuations responsible for marginal transport. Experimentally, $T_{0}$ can be estimated from Raman scattering, where the low-frequency slope of the electronic continuum directly reflects the quasiparticle scattering rate. Defining the slope as $S\equiv\left.\tfrac{dI(\Omega)}{d\Omega}\right|_{\Omega\to 0^{+}}$, one can compare two samples at the same temperature and construct the ratio
$R_{2}(T)=\frac{S_{\rm sample}(T)}{S_{\rm ref}(T)}.$ (60)
Within the Tsallis parametrization $q(T)=1+T/T_{0}$, this ratio admits the inversion formula
$T_{0}(T)=\frac{T}{R_{2}(T)-1},$ (61)
so that experimental Raman slopes immediately yield an estimate of $T_{0}$.
For cuprates, Chang et al. [1] demonstrated that at the critical doping $p_{c}\approx 0.19$, the electronic self-energy evolves from an incoherent regime ($\mathrm{Im}\,\Sigma\sim\text{const.}$) to a coherent Fermi-liquid regime ($\mathrm{Im}\,\Sigma\sim T^{2}$). The Raman $B_{1g}$ low-frequency slope tracks this crossover: it is $\sim 0.5$–1 eV-1 near optimal doping ($p\approx 0.16$) and increases significantly for $p>p_{c}$ in the strange-metal regime. Complementary thermodynamic measurements by Talluri et al. [2] show that the electronic specific heat coefficient $\gamma(T)=C_{\rm el}/T$ is suppressed for $p<p_{c}$, but converges to a linear-in-$T$ form above $p_{c}$, reinforcing the identification of $p_{c}$ as the strange-metal–pseudogap boundary. In iron-based superconductors (Ba(Fe1-xCox)2As2, BFCA), Raman $B_{1g}$ slopes reported by Gallais and Paul [3] and Jost et al. [4] display a similar enhancement near the nematic quantum critical point. Using Eq. (61), slope ratios such as $(n_{\rm dop}=0.04;0.02)$ and $(n_{\rm dop}=0.24;0.07)$ give $T_{0}\simeq 166$–233 K, while for Bi2212 cuprates, comparing $(n_{\rm dop}=0.21;0.19)$ yields $T_{0}\simeq 154$–200 K, consistent across independent datasets.
Taken together, these results show that in both iron-based and cuprate superconductors the extracted Tsallis scale $T_{0}$ clusters around $\mathcal{O}(10^{2})\,\text{K}$. This convergence demonstrates that soft-mode fluctuations of comparable strength underpin marginal scattering, Raman continua, specific heat anomalies, and ultimately high-$T_{c}$ superconductivity across distinct material families. The unified scale $T_{0}$ suggests a universal energy scale for quantum critical fluctuations driving high $T_{c}$ superconductivity, linking Raman, specific heat, and transport anomalies across cuprates and iron-based superconductors.

### v2

@@ Abstract
The reparametrization mode is intrinsically a $0+1$-dimensional object. It governs local energy scrambling and does not carry spatial momentum. Its role in a higher-dimensional metal is therefore limited to providing the local energy-relaxation rate $\Gamma_{Q}$, which follows from the Ohmic spectral density of the Schwarzian sector. All spatial structures, including electric current, momentum conservation, momentum relaxation through umklapp or disorder, and the Lorentz-force contribution to the antisymmetric block $N$, originate from higher-dimensional electronic degrees of freedom. These ingredients are combined by the memory-matrix formalism, which projects local energy dissipation and momentum relaxation onto slow operators. This dimensional separation makes the two relaxation channels robust: the longitudinal resistivity follows $\rho_{xx}\propto\Gamma_{Q}\sim T$, while the Hall angle is governed by the momentum-relaxation rate $\Gamma_{P}\sim T^{2}$, giving $\cot\theta_{H}\propto T^{2}$. The canonical scaling of the Hall angle, $\cot\theta_{H}\propto T^{2}$, observed in optimally doped cuprates and several strange metals, has been widely interpreted as evidence for distinct transport and Hall scattering rates [53]. This behavior, however, is not universal. In electron-doped cuprates, such as $\mathrm{Pr_{2-x}Ce_{x}CuO_{4-\delta}}$ [51], and in layered ruthenates including $\mathrm{Sr_{2}RuO_{4}}$ and $\mathrm{Sr_{3}Ru_{2}O_{7}}$ [52], clear departures from the $T^{2}$ law have been reported. The coexistence of multiple Fermi-surface sheets with distinct orbital characters leads to orbital-dependent scattering lifetimes that evolve differently with temperature and magnetic field. In $\mathrm{Sr_{2}RuO_{4}}$, for instance, the $\alpha$, $\beta$, and $\gamma$ bands derived from different Ru $4d$ orbitals exhibit markedly different coherence scales and gap-like features [52]. Such orbital-selective effects effectively renormalize the temperature dependence of the transport tensors, producing an apparent breakdown of the simple scaling. These observations point to the necessity of introducing a material-dependent cutoff scale $T_{0}$, which encodes band-structure and orbital complexity within a unified transport framework.
@@
$A(\omega;B)\simeq\frac{\kappa_{b}}{C}\frac{|B|\,|\omega|}{T_{0}+T},$ (43)
@@
$\Gamma(0)=\frac{\kappa\pi}{8C}\frac{|B|\,T^{2}}{T_{0}+T},$ (44)
so that at fixed $T$ the longitudinal resistivity scales linearly with field, $\rho_{xx}\propto|B|$. This mechanism captures key experimental observations in quantum-critical strange metals. In BaFe2(As1-xPx)2, near the quantum critical point the scattering rate follows [54]
$\Gamma\approx\sqrt{(k_{B}T)^{2}+(\mu_{B}B)^{2}},$ (45)
demonstrating nearly identical thermal and magnetic energy scales. Similarly, in overdoped La2-xSrxCuO4, scale-invariant linear-in-$B$ resistivity is observed up to $80~\mathrm{T}$ [55], with slope $\beta\sim k_{B}/\mu_{B}$, consistent with Eq. (44). Together, these results are consistent with a Planckian-limited scaling,
$\Gamma(T,B)\propto\max\{k_{B}T,\mu_{B}B\},$ (46)
@@
Acknowledgments- We would like to thank JianXin Lu, Xiao Hu, Sang-Jin Sin, Yan Liu, Matteo Baggioli, Yan-Gang Miao and Kui Jin for helpful discussions. We would also like to express our gratitude to colleagues in the "Eight Immortals Crossing the Sea" WeChat group for their engaging discussions on topics beyond physics. This work was partially supported by NSFC, China (Grant Nos. 12275166 and 12311540141).
@@ Appendix A DC Conductivity of Fermi liquid
We present the standard Kubo-bubble derivation and show its reduction to the semi-classical form used in the main text.
@@ A.2 DC limit and the quasi-particle approximation
where $\partial_{\epsilon}f(\epsilon)$ restricts contributions to energies within a thermal window $\sim T$ around the Fermi level. Here, $A(\mathbf{k},\epsilon)$ is the single particle spectral function, and $v_{x}(\mathbf{k})=\partial\varepsilon_{\mathbf{k}}/\partial k_{x}$ is the group velocity along the $x$-direction. Near the Fermi surface, the spectral function is well approximated by a Lorentzian form
@@
is the single-particle density of states (DOS). Equation (54) is equivalent to the Boltzmann relaxation-time approximation when vertex corrections are negligible, capturing the essential semiclassical physics of momentum relaxation [50]. In the Landau Fermi liquid, the low-energy scattering rate scales as
@@ Appendix B Fermi-liquid limit: Quadratic magnetoresistance
which arises when the magnetic field does not generate new low frequency scattering channels. This includes impurity- or phonon-dominated elastic scattering (field-independent $\tau_{\rm tr}$) and conventional FL inelastic scattering with $\tau_{\rm tr}^{-1}\propto T^{2}$. The expansion is valid in the weak-field limit $\omega_{c}\tau_{\rm tr}\ll 1$ and continuous spectral densities. Departures from this behavior, such as Landau quantization or field-enhanced soft modes, lead to non-FL magnetoresistance, including the linear-in-$|B|$ response discussed in the main text.
@@
Extraction of the Tsallis Scale from Raman Soft Modes
A central quantity in this analysis is the Tsallis scale $T_{0}$, which sets the strength of low-energy soft modes and controls the deviation from Fermi-liquid behavior through $q(T)=1+T/T_{0}$. This scale can be accessed experimentally from the low-frequency Raman response. In strange metals the Raman spectrum is dominated by a broad electronic continuum. For $\omega\ll T$, the low-frequency slope
$S(T)\equiv\left.\partial_{\omega}\mathrm{Im}\,\chi_{R}(\omega)\right|_{\omega\to 0}$
tracks the inverse incoherent relaxation rate  [1, 2],
$S(T)\;\sim\;\frac{\chi_{R}}{\Gamma_{Q}(T)},$
assuming that the same incoherent channel controlling the dc transport also dominates the low-frequency Raman response. Note that $\chi_{R}(\omega)$ is the Raman susceptibility defined as $\chi_{R}(\omega)=-i\int_{0}^{\infty}dt\,e^{i\omega t}\langle[R(t),R(0)]\rangle$, with Raman operator $R=\sum_{\mathbf{k}}\gamma_{\mathbf{k}}\,c^{\dagger}_{\mathbf{k}}c_{\mathbf{k}}$. Within the Tsallis–soft–mode phenomenology introduced above, the Schwarzian soft modes have an Ohmic spectrum $A_{\varepsilon}(\omega)\sim|\omega|/(T_{0}+T)$, which leads to
$\Gamma_{Q}(T)\;\sim\;\frac{T^{2}}{T_{0}+T},$
up to weakly temperature-dependent prefactors. This form yields Fermi-liquid scaling $\Gamma_{Q}\sim T^{2}/T_{0}$ for $T\ll T_{0}$ and Planckian scaling $\Gamma_{Q}\sim T$ for $T\gg T_{0}$. Consequently
$S(T)\;\propto\;\frac{T_{0}+T}{T^{2}}.$
Because the measured Raman intensity satisfies $I(\Omega)\propto[1+n_{B}(\Omega)]\,\mathrm{Im}\,\chi_{R}(\Omega)$, the experimental low–frequency slope $S_{\rm exp}(T)$ is proportional to $S(T)$ up to a known Bose factor. For two samples measured at the same temperature, this factor cancels in the ratio
$R_{2}(T)\;\equiv\;\frac{S_{\rm sample}(T)}{S_{\rm ref}(T)}\;\approx\;\frac{T_{0}^{\rm(sample)}+T}{T_{0}^{\rm(ref)}+T},$
which allows for a controlled extraction of the Tsallis scale $T_{0}$:
$T_{0}^{\rm(sample)}=R_{2}(T)\bigl[T_{0}^{\rm(ref)}+T\bigr]-T.$
The low-energy Raman continuum in underdoped cuprates is naturally connected to the non-quasiparticle dynamics identified in the tomographic-density-of-states analysis of Reber et al. [3]. Their measurements showed that the Fermi-arc spectral weight originates from pair-breaking processes that redistribute spectral weight from the gap edge to the Fermi level rather than from genuine quasiparticle poles. This establishes the antinodal sector as a source of pronounced incoherence.
The restoration of coherent quasiparticles is most clearly observed in the nodal sector. High-resolution ARPES measurements by Chen et al. [4] demonstrated that the strange-metal state in Bi2212 is sharply terminated at a critical doping $p_{c}\simeq 0.19$. For $p<p_{c}$ the nodal excitations remain incoherent with $\mathrm{Im}\,\Sigma\sim\mathrm{const}$, whereas for $p>p_{c}$ they recover the Fermi-liquid form $\mathrm{Im}\,\Sigma\propto T^{2}$. This transition coincides with a pronounced increase of the $B_{1g}$ Raman slope and with the recovery of a temperature-independent electronic specific heat coefficient $\gamma(T)$ [5], providing thermodynamic evidence for the onset of coherence.
In the iron-pnictide Ba(Fe1-xCox)2As2, Raman spectroscopy provides a direct probe of nematic fluctuations. Gallais et al. [6] established that the $B_{1g}$ Raman response is strongly enhanced near the nematic quantum critical composition, reflecting a divergent nematic susceptibility. Kretzschmar et al. [7] mapped the full momentum- and frequency-dependent susceptibility and confirmed that the Raman geometry $B_{1g}$ isolates the $\mathbf{q}=0$ nematic soft mode. Applying the analysis of the ratio to Raman slopes in this critical region yields $T_{0}\simeq 160$$230\,\mathrm{K}$, consistent with the scale extracted in cuprates.
Across both cuprates and iron-based superconductors, Raman spectroscopy, ARPES, and thermodynamic measurements converge on a common Tsallis scale,
$T_{0}\sim 100\,\mathrm{K}.$
This scale reflects the strength of low-energy collective fluctuations, including pseudogap pair-breaking in cuprates and nematic soft modes in iron pnictides, that generate marginal scattering and govern the onset of quasiparticle coherence. The consistency across two distinct material families indicates that $T_{0}$ serves as a unifying measure of the soft-mode physics underlying strange-metal behavior and high-temperature superconductivity.